Enviro Company issues 8%, 10-year bonds with a par value of $260,000 and semiannual interest payments. On the issue date, the annual market rate for these bonds is 10%, which implies a selling price of 87 1⁄2. The straight-line method is used to allocate interest expense.

1. Using the implied selling price of 87 ½, what are the issuer's cash proceeds from issuance of these bonds?
2. What total amount of bond interest expense will be recognized over the life of these bonds?
3. What is the amount of bond interest expense recorded on the first interest payment date?
Answer #1

1) Net Cash Proceeds = 260,000*0.875 = $ 227,500 (Answer)

2). Total bond interest Expenses

Cash interest = 260,000*8%*10 years =208,000

Discount on Bond Issue = 260,000-227,500 = $32,500

Total bond Interest Expenses = 208,000+32,500 = $ 240,500 (Answer)

3) Interest Expenses on First interest Payment :

Cash interest = 260,000*8%*6/12 =10,400

Discount Amortized = 32,500/20 =1,625

Interest Expenses = $12,025 (answer)

Please up vote. Thanks!